# Network-Realized Model Predictive Control Part II: Distributed Constraint Management

### v1

@@ 8 Conclusions
As a contribution to the synthesis of distributed constraint management techniques, we have shown that it is possible to elegantly decouple structural design constraints (which represent the focal point of the distributed control paradigm) from the time-based constraints which are ubiquitous in MPC-oriented literature. By leveraging a promising class of distributed control strategies [9, 10, 11] and by adapting a series of set-theoretical notions [3] to the distributed setting, we obtain a computationally efficient control architecture which provides powerful theoretical guarantees and whose components can be designed fully offline.
$\mathrm{row}_{j}(w_{r\ell}[k+1])=$ $-a_{j\ell}\,\mathrm{row}_{1}(w_{r\ell}[k])+\mathrm{row}_{j+1}(w_{r\ell}[k])+K_{j\ell}\tiny\begin{bmatrix}u_{f}[k]+\beta_{f}[k]\\ x[k]+\zeta[k]+u_{s1}[k]+\beta_{s1}[k]\end{bmatrix},\,\forall\,j\in\{1:n_{r\ell}-1\},$ (A.1a)
$\mathrm{row}_{n_{r\ell}}(w_{r\ell}[k+1])=$ $-a_{n_{r\ell}\ell}\,\mathrm{row}_{1}(w_{r\ell}[k])+K_{n_{r\ell}\ell}\tiny\begin{bmatrix}u_{f}[k]+\beta_{f}[k]\\ x[k]+\zeta[k]+u_{s1}[k]+\beta_{s1}[k]\end{bmatrix},$ (A.1b)
$\mathrm{row}_{\ell}(u_{f}[k])=$ $\ \mathrm{row}_{1}(w_{r\ell}[k]),\,\forall\,k\geq k_{0},\,\ell\in\mathcal{A}_{ui}.$ (A.1c)
$x_{i}[k_{0}+1]=$ $\ C_{xi}\xi_{i}[k_{0}+1]+(\psi_{xi}[k_{0}+1]+\theta_{xi}[k_{0}+1]+\delta_{xi}[k_{0}+1]),$
$\qquad\qquad\qquad\qquad\qquad\qquad\quad\psi_{xi}[k_{0}+1]\in\Psi_{xi1},\,\theta_{xi}[k_{0}+1]\in\Theta_{xi1},\,\delta_{xi}[k_{0}+1]\in\Delta_{xi1},$ (A.2a)
$\mathrm{row}_{j}(u_{fi}[k_{0}+1])=$ $\ \mathrm{row}_{j}(C_{ui})\xi_{i}[k_{0}+1]+(\mathrm{row}_{j}(\psi_{ui}[k_{0}+1])+\mathrm{row}_{j}(\theta_{ui}[k_{0}+1])+\mathrm{row}_{j}(\delta_{ui}[k_{0}+1])),$
$\ \mathrm{row}_{j}(\psi_{ui}[k_{0}+1])\in\left(e_{j}^{\top}\Psi_{ui1}\right),\,\mathrm{row}_{j}(\theta_{ui}[k_{0}+1])\in\left(e_{j}^{\top}\Theta_{ui1}\right),$
$\ \mathrm{row}_{j}(\delta_{ui}[k_{0}+1])\in\left(e_{j}^{\top}\Delta_{ui1}\right),\,\forall\,j\in\{1:n_{ui}\}.$ (A.2b)
@@
$\qquad\qquad\psi_{xi}[k_{0}+1]\in\Psi_{xi1},\,\widetilde{\theta}_{xi}[k_{0}+1]\in\widetilde{\Theta}_{xi1},\,\widetilde{\delta}_{xi}[k_{0}+1]\in\widetilde{\Delta}_{xi1},\,\eta_{xi}[k_{0}+1]\in\mathcal{H}_{xi1},$ (A.3a)
$\mathrm{row}_{j}(u_{fi}[k_{0}+1])=$ $\ \left(\left(\mathrm{row}_{j}(C_{ui})\xi_{i}[k_{0}+1]+\mathrm{row}_{j}\left(\widetilde{\theta}_{ui}[k_{0}+1]\right)\right)+\left(\mathrm{row}_{j}(\psi_{ui}[k_{0}+1])+\mathrm{row}_{j}\left(\widetilde{\delta}_{ui}[k_{0}+1]\right)\right)\right)+$
$\quad\ \ +(-\mathrm{row}_{j}(\eta_{ui}[k_{0}+1])),\,\mathrm{row}_{j}(\psi_{ui}[k_{0}+1])\in\left(e_{j}^{\top}\Psi_{ui1}\right),\,\mathrm{row}_{j}\left(\widetilde{\theta}_{ui}[k_{0}+1]\right)\in\left(e_{j}^{\top}\widetilde{\Theta}_{ui1}\right),$
$\qquad\qquad\mathrm{row}_{j}\left(\widetilde{\delta}_{ui}[k_{0}+1]\right)\in\left(e_{j}^{\top}\widetilde{\Delta}_{ui1}\right),\,\mathrm{row}_{j}(\eta_{ui}[k_{0}+1])\in\left(e_{j}^{\top}\mathcal{H}_{ui1}\right),\forall\,j\in\{1:n_{ui}\}.$ (A.3b)
Appendix A Proofs and Auxiliary Results
We begin by proving the result which enables us to constrain the state variables of the first layer subcontrollers.
Proof of Proposition 5.1
The result follows by straightforward algebraic manipulations of the state-space dynamics given in (8). For the sake of clarity, we write these equations out in an explicit manner and we obtain the identities from (A.1a)-(A.1c), which are located at the top of the next page.
We tackle first the inclusions showcased in point $i)$. Assumption $A2)$ implies the fact that $\mathrm{row}_{1}(w_{r\ell}[k_{0}])\in\mathcal{W}_{\ell 1}$ and, due to Assumption $A1)$, using (34a) and (A.1b) yields the inclusion $\mathrm{row}_{n_{r\ell}}(w_{r\ell}[k_{0}+1])\in\mathcal{W}_{\ell n_{r\ell}}$. Moreover, by employing the same arguments for (34b) and by effecting successive substitutions in (A.1a), we retrieve the fact that $\mathrm{row}_{j}(w_{r\ell}[k_{0}+1])\in\mathcal{W}_{\ell j}$, for all $j\in\{2:n_{r\ell}-1\}$. To get that $\mathrm{row}_{1}(w_{r\ell}[k_{0}+1])\in\mathcal{W}_{\ell 1}$, we simply use Assumption $A3)$ in conjunction with (A.1c). To retrieve point $ii)$, recall Assumption $A2)$ and then use (A.1c) along with (32) to get $u_{fi}[k_{0}]\in{\mathcal{U}}_{fi}$. The desired inclusion follows from Assumption $A1)$, (29) and (31).
Moving on, we prove the result which represents the cornerstone of our constraint management technique.
Proof of Proposition 6.1
The proof is constructive in nature and, to begin with, notice that Assumption A2) enables the local computation of the second layer command pairs $u_{s1i}[k_{0}]\in\mathcal{U}_{s1i}$ and $u_{s2i}[k_{0}]\in\mathcal{U}_{s2i}$, for each $i\in\{1:N\}$, as in Step 3 of Algorithm 1. Since the algebraic conditions from (12) are satisfied by the each of the first layer’s subcontrollers, then there is no need to pass on the values of $u_{s1j}[k_{0}]$, for any $j\in\{1:N\}\backslash\mathcal{N}_{i}$, to the $i^{\text{th}}$ subcontroller from the first layer. Therefore, it follows that by communicating and by applying these partial command signals which originate from the second layer, as indicated in Steps 4a and 4b of Algorithm 1, then the implemented command scheme is equivalent to the one depicted Figure 2. Consequently, the implementation described in Algorithm 1 ensures that the closed-loop dynamics of our control scheme’s first layer are those from (16a)-(16d).
By applying the these commands, the $i^{\text{th}}$ area’s state and its first-layer subcontroller’s outputs at time instant $k=k_{0}+1$ will be given by the identities showcased in (A.2a)-(A.2b), located at the top of this page, where
$\xi_{i}[k_{0}+1]=B_{s1i}u_{s1i}[k_{0}]+B_{s2i}u_{s2i}[k_{0}],\vspace{-3mm}$ (64)
@@ Appendix A Proofs and Auxiliary Results
By employing now the identities from (21a)-(21b) and (24), it follows that the variables described in (A.2a)-(A.2b) can be rewritten as in (A.3a)-(A.3b), also located at the top of this page. However, since $u_{s1i}[k_{0}]$ and $u_{s2i}[k_{0}]$ are obtained by solving (42), it follows that the vectors located on the left-hand side of (64) must satisfy $\xi_{i}[k_{0}+1]\in\Xi_{i1},\,\forall\,i\in\{1:N\}$. By considering the expressions of the sets defined in (43)-(45), it follows by a pair of simple algebraic manipulations that
$\scriptsize\begin{array}[]{l}\left(C_{xi}\xi_{i}[k_{0}+1]+\widetilde{\theta}_{xi}[k_{0}+1]\right)\end{array}\in\mathcal{P}_{1i1},\vspace{-4mm}$
and that
$\scriptsize\begin{array}[]{l}\left(\mathrm{row}_{j}(C_{ui})\xi_{i}[k_{0}+1]+\mathrm{row}_{j}\left(\widetilde{\theta}_{ui}[k_{0}+1]\right)\right)\end{array}\in\mathcal{P}_{2(j+\alpha_{ui})1},\vspace{-3mm}$
@@
for all $i\in\{1:N\}$ along with all $j\in\{1:n_{ui}\}$. Moreover, by using the identities given in (A.3a)-(A.3b) in tandem with the set inclusions from (44)-(45) and with the set definitions introduced in (36) and (37), we get that $x_{i}[k_{0}+1]\in\mathcal{X}_{i}$ and that $\mathrm{row}_{j}(u_{fi}[k_{0}+1])\in\mathcal{W}_{(j+\alpha_{ui})1}$ for all $i\in\{1:N\}$ and all $j\in\{1:n_{ui}\}$.
@@
The first conclusion we are able to draw is the fact that $x[k_{0}+1]\in\mathcal{X}$, by virtue of (26). Now, in order to retrieve the remaining inclusions given in our result’s statement, notice that all of the assumptions associated with Proposition 5.1 hold. Thus, by applying the aforementioned result, we get that $\mathrm{row}_{j}(w_{r\ell}[k_{0}+1])\in\mathcal{W}_{\ell j},$ for all $\,i\in\{1:N\},\,\ell\in\mathcal{A}_{ui}$ and $\,j\in\{1:n_{r\ell}\}$, along with the fact that $u_{i}[k_{0}]\in\mathcal{U}_{i},\,\forall\,i\in\{1:N\}$. Finally, the latter inclusions yield $u[k_{0}]\in\mathcal{U}$, by way of (26).
@@
The central idea, which underpins the entire proof, is to show that all of the problems described in (42) are guaranteed to be feasible at an arbitrary time instant $k_{0}$, and considering some prediction horizons $T_{i}>0$ and $\overline{T}_{i}\geq 0$. Once this is achieved, the proof boils down to a series of induction-based arguments in which Proposition 6.1 is repeatedly applied, in order to ensure the sought-after inclusions from the result’s statement.
@@
We begin by choosing an arbitrary area index $i\in\{1:N\}$ and by investigating its associated area of the network. Due to Assumption A2), we have that (48) holds for some $\rho_{i}\geq 1$. Thus, for any $\widehat{\theta}_{i}\in\widehat{\Theta}_{i\rho_{i}}$, there must exist $\widehat{u}_{i}\in\widehat{\mathcal{U}}_{i\rho_{i}}$ and $\widehat{p}_{i}\in\widehat{\mathcal{P}}_{i\rho_{i}}$ such that $\widehat{\theta}_{i}=-M_{i\rho_{i}}\widehat{u}_{i}+\widehat{p}_{i}$ or, equivalently, such that $\widehat{\theta}_{i}+M_{i\rho_{i}}\widehat{u}_{i}=\widehat{p}_{i}$. Given the definitions introduced in (47b), (47c) and (47e), along with the matrices from (46), the obtained equalities can be rewritten as follows
$\widetilde{\theta}_{it}+\sum_{j=1}^{t}F_{i(j-1)}\tiny\begin{bmatrix}u_{1ij}\\ u_{2ij}\end{bmatrix}\normalsize={p}_{it},\,\forall\,t\in\{1:\rho_{i}\},\vspace{-3mm}$ (65)
having employed the fact that
$\widehat{\theta}_{i}=$ $\ \scriptsize\begin{bmatrix}\widetilde{\theta}_{i1}^{\top}&\dots&\widetilde{\theta}_{i\rho_{i}}^{\top}\end{bmatrix}^{\top},\ \widetilde{\theta}_{it}\in\widetilde{\Theta}_{it},\ t\in\{1:\rho_{i}\},$
$\widehat{u}_{i}=$ $\ \scriptsize\begin{bmatrix}u_{1i1}^{\top}&u_{2i1}^{\top}&\dots&u_{1i\rho_{i}}^{\top}&u_{2i\rho_{i}}^{\top}\end{bmatrix}^{\top},$
$\qquad\qquad\qquad u_{1it}\in\mathcal{U}_{s1i},\ u_{2it}\in\mathcal{U}_{s2i},\ t\in\{1:\rho_{i}\},$
$\widehat{p}_{i}=$ $\ \scriptsize\begin{bmatrix}p_{i1}^{\top}&\dots&p_{i\rho_{i}}^{\top}\end{bmatrix}^{\top},\ p_{it}\in\widetilde{\mathcal{P}}_{it},\ t\in\{1:\rho_{i}\}.$
Now, by exploiting (47a) and (47d), it becomes possible to rewrite (65) in a more explicit manner, which yields
@@
$\left\{\scriptsize\begin{array}[]{l}\widetilde{\theta}_{xit}+C_{xi}\displaystyle\sum_{j=1}^{t}A_{si}^{(j-1)}\scriptsize\begin{bmatrix}B_{s1i}&B_{s2i}\end{bmatrix}\scriptsize\begin{bmatrix}u_{1ij}\\ u_{2ij}\end{bmatrix}={p}_{xit},\\ \widetilde{\theta}_{u\ell t}+\mathrm{row}_{(\ell-\alpha_{ui})}(C_{ui})\displaystyle\sum_{j=1}^{t}A_{si}^{(j-1)}\scriptsize\begin{bmatrix}B_{s1i}&B_{s2i}\end{bmatrix}\scriptsize\begin{bmatrix}u_{1ij}\\ u_{2ij}\end{bmatrix}={p}_{u\ell t},\end{array}\right.$ (66)
for all $t\in\{1:\rho_{i}\}$ and all $\ell\in\mathcal{A}_{ui}$, where $\widetilde{\theta}_{xit}\in\widetilde{\Theta}_{xit}$, $\widetilde{\theta}_{u\ell t}\in\left(e_{(\ell-\alpha_{ui})}^{\top}\widetilde{\Theta}_{uit}\right)$, ${p}_{xit}\in\mathcal{P}_{1it}$ and ${p}_{u\ell t}\in\mathcal{P}_{2\ell t}$.

### v4

@@ 8 Conclusions
As a contribution to the synthesis of distributed constraint management techniques, we have shown that it is possible to elegantly decouple structural design constraints (which represent the focal point of the distributed control paradigm) from the time-based constraints which are ubiquitous in MPC-oriented literature. By leveraging a promising class of distributed control strategies [NRF, part1, aug_sparse] and by adapting a series of set-theoretical notions [STMC] to the distributed setting, we obtain a computationally efficient control architecture which provides powerful theoretical guarantees, and whose components can be designed fully offline.
{ack}
The authors would like to thank both Şerban Sabău and Cristian Oară for the many fruitful discussions which ultimately led to the elaboration of this manuscript.
$\mathrm{row}_{j}(w_{r\ell}[k+1])=$ $-a_{j\ell}\,\mathrm{row}_{1}(w_{r\ell}[k])+\mathrm{row}_{j+1}(w_{r\ell}[k])+K_{j\ell}\tiny\begin{bmatrix}u_{f}[k]+\beta_{f}[k]\\ x[k]+\zeta[k]+u_{s1}[k]+\beta_{s1}[k]\end{bmatrix},\,\forall\,j\in\{1:n_{r\ell}-1\},$ (B.1a)
$\hskip-2.84526pt\mathrm{row}_{n_{r\ell}}(w_{r\ell}[k+1])=$ $-a_{n_{r\ell}\ell}\,\mathrm{row}_{1}(w_{r\ell}[k])+K_{n_{r\ell}\ell}\tiny\begin{bmatrix}u_{f}[k]+\beta_{f}[k]\\ x[k]+\zeta[k]+u_{s1}[k]+\beta_{s1}[k]\end{bmatrix}$, (B.1b)
$\mathrm{row}_{\ell}(u_{f}[k])=$ $\ \mathrm{row}_{1}(w_{r\ell}[k]),\,\forall\,k\geq k_{0},\,\ell\in\mathcal{A}_{ui}.$ (B.1c)
$x_{i}[k_{0}+1]=$ $\ C_{xi}\xi_{i}[k_{0}+1]+(\psi_{xi}[k_{0}+1]+\theta_{xi}[k_{0}+1]+\delta_{xi}[k_{0}+1]),\,\psi_{xi}[k_{0}+1]\in\Psi_{xi1},\,\theta_{xi}[k_{0}+1]\in\Theta_{xi1},\,\delta_{xi}[k_{0}+1]\in\Delta_{xi1},$ (B.2a)
$\mathrm{row}_{j}(u_{fi}[k_{0}+1])=$ $\ \mathrm{row}_{j}(C_{ui})\xi_{i}[k_{0}+1]+(\mathrm{row}_{j}(\psi_{ui}[k_{0}+1])+\mathrm{row}_{j}(\theta_{ui}[k_{0}+1])+\mathrm{row}_{j}(\delta_{ui}[k_{0}+1])),\,\mathrm{row}_{j}(\psi_{ui}[k_{0}+1])\in\left(e_{j}^{\top}\Psi_{ui1}\right),$
$\ \mathrm{row}_{j}(\theta_{ui}[k_{0}+1])\in\left(e_{j}^{\top}\Theta_{ui1}\right),\,\mathrm{row}_{j}(\delta_{ui}[k_{0}+1])\in\left(e_{j}^{\top}\Delta_{ui1}\right),\,\forall\,j\in\{1:n_{ui}\}.$ (B.2b)
Appendix A Summary of NRF-based Control
We include here several key definitions and properties associated with the class of control laws from [NRF].
Fundamentally, the concept of an NRF-based representation arises when attempting to represent the TFM of a centralised controller, denoted henceforth as $\mathbf{K}(z)$ and assumed to be of dimension $n_{u}\times n_{x}$, via the expression
$\mathbf{K}(z)=(I-\mathbf{\Phi}(z))^{-1}\mathbf{\Gamma}(z),\vskip-2.84526pt$
for which $\mathrm{det}(I-\mathbf{\Phi}(z))\not\equiv 0$ and $\mathrm{elm}_{ii}({\Phi}(z))\equiv 0$, for all $i\in\{1:n_{u}\}$. Thus, the TFM pair $(\mathbf{\Phi}(z),\mathbf{\Gamma}(z))$ allows for
the computation of the signal $u[k]$ in (1a)-(1c) as follows
$u[k]=\mathbf{\Phi}(z)\star u[k]+\mathbf{\Gamma}(z)\star x[k].\vskip-8.53581pt$ (58)
Since the diagonal elements of $\mathbf{\Phi}(z)$ are all identically zero this enables a distinct feedback-feedforward expression of the resulting control laws, in which an arbitrary command signal $u_{i}[k]$ may be computed via the use of a feedback component $x[k]$ alongside the feedforward of the command signals $u_{j}[k]$ belonging to other nodes in the network ($j\neq i$). It is precisely this feature which, when expressed via (58), yields the identities from (12).
In the NRF framework, the problem of imposing sparsity on $\mathbf{K}(z)$ for a distributed implementation (a highly restrictive objective, see [NRF] and [aug_sparse]) is relaxed into ensuring sparsity for its NRF pair $(\mathbf{\Phi}(z),\mathbf{\Gamma}(z))$. The latter goal is far more tractable, due to the manner in which the pair may be computed. To this end, we now introduce the notion of Doubly Coprime Factorisation (DCF) over a set $\mathbb{C}_{g}$, which can be employed as a design parameter.
Remark A.1.
The choice of $\mathbb{C}_{g}$ underlines several key properties of the resulting NRF-based control laws. For example, choosing  $\mathbb{C}_{g}\subseteq\{z\in\mathbb{C}\,|\,|z|\hskip-1.42262pt<\hskip-1.42262pt1\}$ ensures that the effect of bounded exogenous disturbance on the signals arising in closed-loop configuration remains bounded.
With respect to the realisation from (1a)-(1c), we define $\mathbf{G}_{u}(z):=(zI_{n_{x}}-A)^{-1}B_{u}$ and we say that any collection of eight TFMs $(\mathbf{N}(z),$ $\mathbf{M}(z),\mathbf{X}(z),\mathbf{Y}(z),\widetilde{\mathbf{N}}(z),$ $\widetilde{\mathbf{M}}(z),\widetilde{\mathbf{X}}(z),\widetilde{\mathbf{Y}}(z))$ is a DCF over $\mathbb{C}_{g}$ of $\mathbf{G}_{u}(z)$ if:
  1. a)
    $\mathbf{M}(z)$ along with $\widetilde{\mathbf{M}}(z)$ are invertible and they satisfy $\mathbf{G}_{u}(z)=\mathbf{N}(z)\mathbf{M}^{-1}(z)=\widetilde{\mathbf{M}}^{-1}(z)\widetilde{\mathbf{N}}(z)$;
  2. b)
    all eight TFMs are proper (taking $|z|\rightarrow\infty$ results in a matrix having only finite entries) and have no poles outside of $\mathbb{C}_{g}$;
  3. c)
    the eight TFMs satisfy the (Bézout-like) identity
[…]
Using just such a DCF over $\mathbb{C}_{g}$, it is possible to obtain a proper (and, thus, implementable) NRF pair for a controller of the system in (1a)-(1c) as follows
[…]
for a TFM $\mathbf{Q}(z)$ that is proper, that has no poles outside of $\mathbb{C}_{g}$, and which ensures that $\widetilde{\mathbf{Y}}_{\mathbf{Q}}(z)$ and $\widetilde{\mathbf{Y}}_{\mathbf{Q}}^{\mathrm{diag}}(z)$ have proper inverses. For systems of type (1a)-(1c), the DCF over $\mathbb{C}_{g}$ can always be chosen so that such a $\mathbf{Q}(z)$ exists.
One of the main features of this representation is the fact that, by computing the NRF pair as above, its two terms inherit the sparsity patterns of $\widetilde{\mathbf{Y}}_{\mathbf{Q}}(z)$ and $\widetilde{\mathbf{X}}_{\mathbf{Q}}(z)$. Indeed, since the two TFMs are affine terms of $\mathbf{Q}(z)$, ensuring desired sparsity patterns on these expressions is a particularly tractable problem (see [aug_sparse], for example). A further benefit of employing this control formalism is the fact that all closed-loop maps from exogenous disturbance to the interconnection’s internal signals (recall Remark A.1) are similarly simple expressions of the $\mathbf{Q}(z)$ parameter. For example, the TFM $\mathbf{F}_{\mathbf{Q}}(z)\scriptsize\begin{bmatrix}I_{(n_{x}+n_{u})}&O\end{bmatrix}^{\top}$ from (14) can be expressed as follows
$\mathbf{F}_{\mathbf{Q}}(z)\scriptsize\begin{bmatrix}I_{(n_{x}+n_{u})}\\ O\end{bmatrix}=\left[\begin{array}[]{cc}\mathbf{N}(z)\widetilde{\mathbf{X}}_{\mathbf{Q}}(z)&\mathbf{N}(z)\widetilde{\mathbf{Y}}_{\mathbf{Q}}(z)\\ \mathbf{M}(z)\widetilde{\mathbf{X}}_{\mathbf{Q}}(z)&\mathbf{M}(z)\widetilde{\mathbf{Y}}_{\mathbf{Q}}(z)-I_{n_{u}}\end{array}\right],\vskip-8.53581pt$
with the other TFMs from (14) being given explicitly in the main result of our companion paper [part1]. Consequently, systemic norm-based procedures (such as the ones proposed in [aug_sparse]) can be used to manipulate these amenable expressions, with the aim of decoupling the areas of the network from (1a)-(1c), as per Remark 4.4.
Appendix B Proofs and Auxiliary Results
We begin by proving the result, which enables us to constrain the state variables of the first-layer subcontrollers.
Proof of Proposition 5.1
The result follows by straightforward algebraic manipulations of the state-space dynamics given in (10). For the sake of clarity, we write these equations out in an explicit manner and we obtain the identities from (B.1a)-(B.1c), which are located at the top of this page.
First, we tackle the inclusions showcased in point i). Assumption A2) implies the fact that $\mathrm{row}_{1}(w_{r\ell}[k_{0}])\in\mathcal{W}_{\ell 1}$ and, due to Assumption A1), using (34a) and (B.1b) yields the inclusion $\mathrm{row}_{n_{r\ell}}(w_{r\ell}[k_{0}+1])\in\mathcal{W}_{\ell n_{r\ell}}$. Moreover, by employing the same arguments for (34b) and by effecting successive substitutions in (B.1a), we retrieve the fact that $\mathrm{row}_{j}(w_{r\ell}[k_{0}+1])\in\mathcal{W}_{\ell j}$, for all $j\in\{2:n_{r\ell}-1\}$. To get that $\mathrm{row}_{1}(w_{r\ell}[k_{0}+1])\in\mathcal{W}_{\ell 1}$, we simply use Assumption A3) in conjunction with (B.1c). To retrieve point ii), recall Assumption A2) and then use (B.1c) along with (32) to get $u_{fi}[k_{0}]\in{\mathcal{U}}_{fi}$. The desired inclusion follows from Assumption A1), (29) and (31).
Moving on, we now prove the result which represents the cornerstone of our constraint management technique.
Proof of Proposition 6.1
The proof is constructive in nature and, to begin with, notice that Assumption A2) enables the local computation of the second-layer command pairs $u_{s1i}[k_{0}]\in\mathcal{U}_{s1i}$ and $u_{s2i}[k_{0}]\in\mathcal{U}_{s2i}$, for each $i\in\{1:N\}$, as in Step 3 of Algorithm 1. Since the algebraic conditions from (13) are satisfied by each of the first layer’s subcontrollers, then there is no need to pass on the values of $u_{s1j}[k_{0}]$, for any $j\in\{1:N\}\setminus\,\mathcal{N}_{i}$, to the $i^{\text{th}}$ subcontroller from the first layer. Therefore, it follows that by communicating and by applying these partial command signals which originate from the second layer, as indicated in Steps 4a and 4b of Algorithm 1, then the implemented command scheme is equivalent to the one depicted in Figure 2. Consequently, the implementation described in Algorithm 1 ensures that the closed-loop dynamics of our control scheme’s first layer are those in (16a)-(16d).
By applying the above-mentioned commands, the $i^{\text{th}}$ area’s state and its first-layer subcontroller’s outputs at time $k=k_{0}+1$ will be given by the identities showcased in (B.2a)-(B.2b), located at the top of this page, where
$\xi_{i}[k_{0}+1]=B_{s1i}u_{s1i}[k_{0}]+B_{s2i}u_{s2i}[k_{0}],\vskip-2.84526pt$ (61)
@@ Appendix B Proofs and Auxiliary Results
$\hskip 16.38895pt\hskip 16.38895pt\hskip 16.38895pt\hskip 16.38895pt\hskip 16.38895pt\hskip 16.38895pt\psi_{xi}[k_{0}+1]\in\Psi_{xi1},\,\widetilde{\theta}_{xi}[k_{0}+1]\in\widetilde{\Theta}_{xi1},\,\widetilde{\delta}_{xi}[k_{0}+1]\in\widetilde{\Delta}_{xi1},\,\eta_{xi}[k_{0}+1]\in\mathcal{H}_{xi1},$ (B.4a)
$\mathrm{row}_{j}(u_{fi}[k_{0}+1])=$ $\left(\left(\mathrm{row}_{j}(C_{ui})\xi_{i}[k_{0}+1]+\mathrm{row}_{j}\left(\widetilde{\theta}_{ui}[k_{0}+1]\right)\right)\hskip-1.42262pt+\hskip-1.42262pt\left(\mathrm{row}_{j}(\psi_{ui}[k_{0}+1])\hskip-1.42262pt+\hskip-1.42262pt\mathrm{row}_{j}\left(\widetilde{\delta}_{ui}[k_{0}+1]\right)\right)\right)+(-\mathrm{row}_{j}(\eta_{ui}[k_{0}+1])),$
$\qquad\ \mathrm{row}_{j}(\psi_{ui}[k_{0}+1])\in\left(e_{j}^{\top}\Psi_{ui1}\right),\,\mathrm{row}_{j}\left(\widetilde{\theta}_{ui}[k_{0}+1]\right)\in\left(e_{j}^{\top}\widetilde{\Theta}_{ui1}\right),\,\mathrm{row}_{j}\left(\widetilde{\delta}_{ui}[k_{0}+1]\right)\in\left(e_{j}^{\top}\widetilde{\Delta}_{ui1}\right),$
$\qquad\ \mathrm{row}_{j}(\eta_{ui}[k_{0}+1])\in\left(e_{j}^{\top}\mathcal{H}_{ui1}\right),\forall\,j\in\{1:n_{ui}\}.$ (B.4b)
@@
By employing now the identities from (21a)-(21b) and (24), it follows that the variables described in (B.2a)-(B.2b) can be rewritten as in (B.4a)-(B.4b), located at the top of the next page. However, since $u_{s1i}[k_{0}]$ and $u_{s2i}[k_{0}]$ are obtained by solving (8), it follows that the vectors located on the left-hand side of (61) must satisfy
$\xi_{i}[k_{0}+1]\in\Xi_{i1},\,\forall\,i\in\{1:N\}.\vskip-8.53581pt$
By considering the expressions of the sets defined in (42)-(44), it follows by simple algebraic manipulations that
[…]
@@
for all $i\in\{1:N\}$, along with all $j\in\{1:n_{ui}\}$. Moreover, by using the identities which form (B.4a)-(B.4b) in tandem with the set inclusions from (42)-(43) and with the set definitions introduced in (36) and (37), we get that $x_{i}[k_{0}+1]\in\mathcal{X}_{i}$ and that $\mathrm{row}_{j}(u_{fi}[k_{0}+1])\in\mathcal{W}_{(j+\alpha_{ui})1}$ for all $i\in\{1:N\}$ and all $j\in\{1:n_{ui}\}$.
@@
The first conclusion we are able to draw, at this point, is the fact that $x[k_{0}+1]\in\mathcal{X}$, by virtue of (26). Now, in order to retrieve the remaining inclusions given in our result’s statement, notice that all of the assumptions associated with Proposition 5.1 hold. Thus, by applying said result, we get that $\mathrm{row}_{j}(w_{r\ell}[k_{0}+1])\in\mathcal{W}_{\ell j},$ for all $i\in\{1:N\},\,\ell\in\mathcal{A}_{ui}$ and $j\in\{1:n_{r\ell}\}$, along with the fact that $u_{i}[k_{0}]\in\mathcal{U}_{i},\,\forall\,i\in\{1:N\}$. Finally, the latter inclusions yield $u[k_{0}]\in\mathcal{U}$, by way of (26).
@@
The central idea, which underpins the entire proof, is to show that all of the problems described in (8) are guaranteed to be feasible at an arbitrary time instant $k_{0}$, and considering some prediction horizons $T_{i}>0$ and $\overline{T}_{i}\geq 0$. Once this is achieved, the proof boils down to a series of induction-based arguments in which Proposition 6.1 is repeatedly applied, with the aim of ensuring the sought-after inclusions from the result’s statement.
@@
We begin by choosing an arbitrary area index $i\in\{1:N\}$, and by investigating its associated area of the network. Due to Assumption A2), we have the fact that (46) holds for some $\rho_{i}\geq 1$. Thus, for any $\widetilde{\theta}_{it}\in\widetilde{\Theta}_{it}$ and any $t\in\{1:\rho_{i}\}$, there must exist $p_{it}\in\widetilde{\mathcal{P}}_{it}$, $u_{1it}\in\mathcal{U}_{s1i}$ along with $u_{2it}\in\mathcal{U}_{s2i}$ such that the following identities hold
@@
$\scriptsize\begin{array}[]{l}\widetilde{\theta}_{it}=\left(\begin{bmatrix}-C_{xi}\\ -C_{ui}\end{bmatrix}\displaystyle\sum_{j=1}^{t}A_{si}^{j-1}(B_{s1i}u_{1ij}+B_{s2i}u_{1ij})\right)+{p}_{it}.\end{array}\vskip-8.53581pt\normalsize\hskip-5.69054pt$ (63)
Now, by exploiting (45a) and (45b), it becomes possible to rewrite (63) in a more explicit manner, which yields
$\left\{\scriptsize\begin{array}[]{l}\widetilde{\theta}_{xit}+C_{xi}\displaystyle\sum_{j=1}^{t}A_{si}^{(j-1)}\scriptsize\begin{bmatrix}B_{s1i}&B_{s2i}\end{bmatrix}\hskip-2.84526pt\scriptsize\begin{bmatrix}u_{1ij}\\ u_{2ij}\end{bmatrix}\hskip-2.84526pt={p}_{xit},\\ \widetilde{\theta}_{u\ell t}+\mathrm{row}_{(\ell-\alpha_{ui})}(C_{ui})\displaystyle\sum_{j=1}^{t}A_{si}^{(j-1)}\scriptsize\begin{bmatrix}B_{s1i}&B_{s2i}\end{bmatrix}\hskip-2.84526pt\scriptsize\begin{bmatrix}u_{1ij}\\ u_{2ij}\end{bmatrix}\hskip-2.84526pt={p}_{u\ell t},\end{array}\right$. (64)